OpenStudy (anonymous):

Given the function g(x) = 8x - 2, find g(-2). Answer 14 -18 -14 0.5

OpenStudy (anonymous):

g(x) = 8x - 2, find g(-2). When they notate g(-2) it means to substitute x with -2 So doing that you will get, g(-2) = 8(-2) - 2 = -16-2 =-18

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Np!